Subject: APO 3.5m users committee minutes from meeting on 5/3/10
From: Suzanne Hawley
Submitted: Tue, 4 May 2010 13:08:27 -0700 (PDT)
Message number: 1181
(previous: 1180,
next: 1182
up: Index)
APO 3.5-m Users Committee Phonecon, 5/03/10
Attending: Suzanne Hawley, Mark Klaene, Michael Strauss, Bill Ketzeback,
John Bally, Jon Fulbright, Jon Holtzman, David Wehlan (for Remy Indebetouw),
Scott Anderson, Bruce Gillespie
Absent: Al Harper
**********************************
User feedback, comments from institutional representatives:
Chicago (Al Harper) - no report.
New Mexico State (Jon Holtzman) - Jon had nothing new to report.
Washington (Scott Anderson) - Scott had nothing new to report.
Colorado (John Bally) - John had nothing new to report.
Princeton (Michael Strauss) - Michael said that everything was fine.
Johns Hopkins (Jon Fulbright) - Jon said that he had sent Suzanne
comments on the proposed RITMOS instrument.
Virginia (David Wehlan) - David had nothing new to report.
**********************************
Discussion of telescope/instruments report - Klaene:
Mark mentioned that the summer shutdown dates have been set, 26 July through
13 August. He also highlighted recent work on refining the telescope and
primary mirror motion control systems. We discussed at some length the
question of improving high-altitude tracking from a cost-benefit standpoint.
Suzanne pointed out that an increasing number of programs are using Agile
to monitor exo-planet transits, and tracking them through the near-zenith
is becoming a priority. This led to the following action item:
===> ACTION: Users Committee members should poll their users to find out
how important it is to their science programs to track targets within
5 degrees of zenith.
Mark also pointed to the newly commissioned ability to on-chip guide with
Agile, and the plan to migrate the Agile guiding software from TUI to the Hub.
Regarding the newly revived SPIcam, users should watch for any performance
anomalies and report them to the observatory, now that the instrument is
back on the air. Mark also recapped the status and plans for the new
NIC-FPS fanout board, and the spare telescope drive assembly.
*************************
3.5-m Telescope, Instruments, and CIF Projects Highlights, 4/2/10 through 4/28/10
0) Overview
Despite last week's 6-inch snowfall, spring is arriving, and some of the
facility work that has been put on hold for the winter is starting. We
are expecting a busy summer with a student group in July along with the
2.5-m shutdown and one retirement (Norm Blythe, who has handled all of
our LN2 usage for the past 7 years). The 3.5-m shutdown dates have been set:
start of the shutdown is July 26 and we will resume science on August 14,
if the weather cooperates.
1) Telescope
Pointing and tracking behavior is still quite good with rare minor issues
with a drive resonance on the azimuth axis; Fritz and Nick are working this.
The problem doesn't seem to have a significant effect on telescope performance,
but it needs to be corrected. An electrical noise problem with the primary
mirror support system was identified and corrected; this appears to have
eliminated the crosstalk and oscillations we have been seeing on some sectors.
2) Instruments
DIS, TripleSpec, NIC-FPS, the Echelle, and Agile ran with no problems, although
the DIS blue camera temperature is still not quite as stable as we would like.
The DIS slitviewer ICC was replaced with a modern machine.
Agile on-chip guiding has completed its testing and is now fully implemented
in TUI. The guiding coefficients are the same as used for the rest of our
guiders, but they can be modified by the user if the default values are not
ideal for specific programs. Users should read the documentation in the
latest TUI release and for Agile before using this feature. The current
implementation is intended to to be a stopgap until we can complete modifications
to the Hub software, where the other guiders are commanded. Note that because
Agile guiding is implemented through TUI and not through the Hub, only one
user can actively use it at a time. Therefore, users must coordinate with
the Observing Specialist and among themselves if using multiple log-ins.
SPIcam has been inoperable due to a failed cryotiger compressor for some time.
We went through the four spare compressors that we had on site, and each of
them had a problem. A new compressor just arrived and has been installed,
and the instrument is now cold. Although we are currently monitoring
temperature stability, chip contamination, and bias levels, SPIcam has been
returned to service for science programs. Any SPIcam performance problems
that are noticed by users should immediately be brought to the attention
of the Observing Specialist and the Observatory.
The new NIC-FPS fanout board testing at CU has been slowed due to problems
with the spare Leach controller. It was decided to investigate and try to
fix these problems and continue with lab testing with the engineering-grade
chip before attempting installing the board on the instrument itself. This
will delay the installation of the new fanout board at the Observatory.
There was still no progress reported on the NIC-FPS etalon at CU.
3) CIF projects
The spare telescope drive box machining and fabrication continues; some
parts are assembled and additional spare components have been delivered.
Nick has prepared a list of mechanical and optical improvement projects
with preliminary designs and budgetary estimates. The list is being vetted
by engineering staff and management to see which items are high priority,
and which could be implemented this year. One of these projects is the
investigation of the high-altitude tracking limit.
**********************************
Agile on-chip guiding - Hawley:
Suzanne added that the recent testing of the on-chip guiding with Agile shows
a 1 to 2 pixel stability over long periods, and feedback from other users
is welcome. Kudos to Russell Owen for getting this capability implemented
in a relatively short time.
**********************************
SPIcam back in service - Klaene:
Aside from the discussion given in the telescope and instruments report,
Mark said that we are purchasing a new spare cryotiger for the instrument
so that we should not have extended down time if the currently installed
cryotiger were to fail.
**********************************
Wyoming IR camera - Hawley:
Suzanne reminded the Users Committee that this proposed instrument has been
in discussion for several years. It would provide IR imaging over ~13
arcminutes, and ultimately provide multi-object spectroscopy at R = ~3000 in
J, H, or K using a microshutter array. The spectroscopic option has a heritage
traceable to the GSFC RIVMOS instrument. One serious issue that needs
exploration is that the current (baffled) FOV at the NA2 port is only 6-7
arcminutes, and the NA2 guider will also partially occult this instrument.
We will conduct an Instrument Review, possibly this summer, to evaluate and
determine the necessary interface and operational aspects of the instrument.
Suzanne would like one or more of the Users Committee members to participate
in this review. If the review is successfully completed, then we will look
toward a visiting instrument run, likely block scheduled in Q4 2010 or beyond.
The instrument PI, Mike Pierce, has provided a white paper that has been sent
to the Users Committee members; this can be distributed to users at each
institution. He has also invited people to contact him directly to discuss
the instrument.
===> ACTION: Users Committee members should forward the NIIS white paper
to their users, and potential users should send any comments or questions
to Suzanne Hawley.
**********************************
Summer Engineering Plans - Klaene:
Mark listed the projects and activities planned for this summer:
o secondary and tertiary mirrors will be realuminized, and their
actuator systems will be serviced;
o primary mirror will be washed (it will be realuminized at Kitt Peak next summer);
o new circuit boards for the direct-drive motor controllers will be installed;
o instruments (except NIC-FPS and echelle) will receive their annual servicing; and
o enclosure will receive its routine maintenance.
NIC-FPS servicing will await the installation of the new fanout board. Echelle
was extensively serviced last year and is performing well. It will be slated
for additional work in the coming year as the upgrade project moves forward.
Suzanne pointed out that for the first time in many years, there is no
major upgrade planned for the telescope or instruments this summer. This reflects
the attainment of a certain level of maturity of the telescope systems hardware,
and accounts for the shorter-than-usual shutdown period. Hence, there will be
more observing time available this summer than we are accustomed to, although
how much of it will be usable during the monsoon season remains to be seen.
**********************************
ACTION ITEMS:
[open from previous months]:
None
[new actions from this meeting]:
===> ACTION: Users Committee members should poll their users to find out
how important it is to their science programs to track targets within
5 degrees of zenith.
===> ACTION: Users Committee members should forward the NIIS white paper
to their users, and potential users should send any comments or questions
to Suzanne Hawley.
**********************************
Next meeting: The next Users Committee phonecon will be on Monday, 14 June,
at 8:30 AM Pacific Time. The agenda and other materials will be sent
to the committee members during the preceding week.
APO APO APO APO APO Apache Point Observatory 3.5m APO APO APO
APO
APO This is message 1181 in the apo35-general archive. You can find
APO the archive on http://www.astro.princeton.edu/APO/apo35-general/INDEX.html
APO To join/leave the list, send mail to apo35-request@astro.princeton.edu
APO To post a message, mail it to apo35-general@astro.princeton.edu
APO
APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO